# Primes and polygonal numbers

### v1

@@ Abstract.
For all integers $r,s>2$, there are infinitely many primes which can be written as the sum of an $r$-gonal number and an $s$-gonal number. In general, any linear combination of an $r$-gonal number and an $s$-gonal number with coprime positive integer coefficients produces infinitely many primes. However, unless $r=s=4$, such primes can not be described by finitely many congruences.
@@ 1. Introduction and the main result
A prime is a sum of two squares if and only if it is of the form $14k+$ 1 (see [8]). Since there are infinitely many primes of the form $14k+$ 1, it follows that there are infinitely many primes which are sum of two squares. It is natural to wonder whether the same holds for triangular numbers or more generally, whether for all integers $r,s\geq 3$, there are infinitely many primes which can be represented as the sum of an $r$-gonal number and an $s$-gonal number, where the $m$-th $r$-gonal number is defined by (see Chapter 2 in [2])
@@
Even more generally, one may ask whether for each pair of integers $r,s\geq 3$ and for each pair of mutually coprime positive integers $\mu$ and $\nu$, there are infinitely many primes of the form
$\mu T(r,m)+\nu T(s,n).$
We answer this in affirmative.
@@
For sufficiently large $N$, for each pair of integers $r,s\geq 3$ and for each pair of mutually coprime positive integers $\mu$ and $\nu$, the number of primes of the form
$\mu T(r,m)+\nu T(s,n)$
@@
By an admissible arithmetic progression, we mean an integer arithmetic progression that has a pair of mutually coprime terms.
@@
Unless $r=s=4$, for each pair of mutually coprime positive integers $\mu$ and $\nu$, the primes of the form
$\mu T(r,m)+\nu T(s,n),$
can not be described as the primes contained in a set of admissible arithmetic progressions.
In other words, one can only describe the primes of the form
$ax^{2}+by^{2}$
by a finite set of congruences for certain $a,b\in\mathbb{N}$ (see [3]). However, Theorem 1 together with the Prime Number Theorem for arithmetic progressions (Corollary 11.20 in [7]) imply that for any pair of $a,b\in\mathbb{N}$, if we replace either $x^{2}$ or $y^{2}$ above by another polygonal number, the resulting linear combination of polygonal numbers misses infinitely many primes in every admissible arithmetic progression.
In particular, the sequence $\{4k+1\}_{k\in\mathbb{N}}$ or its subprogressions are the only admissible arithmetic progressions such that each prime in those progressions is a sum of two polygonal numbers.
It follows from Theorem 1 and Chebyshev’s theorem (Corollary 2.6 in [7]) that the set of the primes, which are sum of two polygonal numbers (or more generally, linear combination of two polygonal numbers with coprime positive integer coefficients) at least one of which is not a square, does not have positive relative upper density with respect to the set of all primes. So, Green-Tao Theorem [4] which guarantees the existence of arbitrarily long arithmetic progressions among the subsets of the primes with positive relative upper densities, is not applicable here either.
From Theorem 1, for $\mu=\nu=1$ and $r=s=3$, we obtain the following corollary which answers the Open Problem 1 in [1].
@@
Since a number $n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares (see [1]), we have
@@
There are infinitely many primes $p$ such that $14p+$ 1 is a sum of two squares.
It is natural to wonder whether there are infinitely many primes $p$ such that both $p$ and $14p+$ 1 are sums of two squares or more generally, whether both the residue classes $1$ and $3$ modulo $4$ contain infinitely many primes $p$ such that $14p+$ 1 is a sum of two squares. Even more generally, one may ask:
@@
Does every admissible arithmatic progression contain infinitely many primes which are sum of two triangular numbers?
2. Proof of the main result
@@
A Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g$ be of an irreducible polynomial of degree $2$ in $\mathbb{Z}[x,y]$ such that $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly independent and $P(x,y)$ represents arbitrarily large odd numbers for $x,y\in\mathbb{N}$. Then
@@
Since for all $i\in\{1,2,3\}$, the coefficients of $x^{2}$ and $y^{2}$ in $P_{i}(x,y)$ are positive and even, whereas the coefficient of $xy$ is zero, it follows that if any of the remaining coefficients of $P_{i}$ are odd, then $P_{i}(x,y)$ produces infinitely many odd numbers for $x,y\in\mathbb{N}$. In particular, if the greatest common divisors of the coefficients of $P_{1}$, $P_{2}$, $P_{3}$ are 1, then each of these polynomials represents arbitrarily large odd numbers. In the following we compute the gcd $G$ of the coefficients of $P_{i}$ for $i\in\{1,2,3\}$.
@@
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{1}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{\mu r,\nu s\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise. }\end{cases}$
@@
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{2}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\nu\notin 2\mathbb{Z}$}\\ \\ \asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise. }\end{cases}$
@@
$\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{2}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1\hskip 5.0pt\begin{cases}\asymp\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\mu+\nu\notin 2\mathbb{Z}$}\\ \\ \asymp\dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,s,\mu+\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ \leq 1&\text{otherwise. }\end{cases}$

### v3

@@ Abstract.
For all integers $r,s>2$, each linear combination of an $r$-gonal number and an $s$-gonal number with coprime positive integer coefficients produces infinitely many primes. However, for any given pair of coprime positive integer coefficients of the polygonal numbers, the sum of the reciprocals of such primes diverges if and only if $r=s=4$.
@@ 1. Introduction and the main result
A prime is a sum of two squares if and only if it is of the form $14k+$ 1. Since there are infinitely many primes of the form $14k+$ 1, it follows that there are infinitely many primes which are sum of two squares. Since the sums of couples of triangular numbers are intimately related with the sums of couples of squares, viz. $n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares, it is natural to expect that the primes which are sums of two triangular numbers will also have an equally appealing description in terms of some arithmetic progressions. However, it turns out not to be the case. In fact, the sum of the reciprocals of such primes converges. One may still wonder whether there are infinitely many such primes or more generally, whether for a given pair of integers $r,s>2$, there are infinitely many primes which are sum of an $r$-gonal number and an $s$-gonal number, where the $m$-th $r$-gonal number is defined by (see [2])
@@
Even more generally, one may ask whether for each pair of integers $r,s>2$ and for every pair of mutually coprime positive integers $a$ and $b$, there are infinitely many primes of the form $aT(r,m)+bT(s,n).$ We answer this in the affirmative.
@@
Given a pair of integers $r,s>2$ and a pair of mutually coprime positive integers $a$ and $b$, for all $m,n\in\mathbb{N}$, the number of primes of the form
$aT(r,m)+bT(s,n)$ (1.2)
@@
From the above theorem, via Abel’s summation formula, we obtain
@@
Given $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, the sum of the reciprocals of the primes of the form $aT(r,m)+bT(s,n)$ diverges if and only if $r=s=4$.
In particular, Theorem 1 solves an open problem in [1], which asks how many primes are sums of two triangular numbers.
@@
In fact, there are infinitely many such primes in each of the residue classes ±$\pm 1$ modulo $4$.
Theorem 2.
Given two mutually coprime positive integers $a$ and $b$, the number of the primes in each of the residue classes ±$\pm 1$ modulo $4$, which are less than $N$ and which are linear combinations of two odd-sided polygonal numbers with coefficients $a$ and $b$ is
$\asymp\dfrac{N}{(\log N)^{3/2}}.$
Since a number $n$ is a sum of two triangular numbers if and only if $14n+$ 1 is a sum of two squares (see [1]) and since a prime is a sum of two squares if and only if it is of the form $14k+$ 1 (see [8]), we have
@@
The number of primes $p\leq N$ such that both $p$ and $14p+$ 1 are a sums of two squares is
$\asymp\dfrac{N}{(\log N)^{3/2}}.$
By an admissible arithmetic progression, we mean an integer arithmetic progression that has a pair of mutually coprime terms. Now, looking at Theorem 2, one may wonder whether every admissible arithmetic progression contains infinitely many primes which are sums of two triangular numbers. However, that does not hold in general! For example, if $n$ is congruent to $5$ or $8$ modulo 9, then $14n+$ 1 is not a sum of two squares. So, none of the arithmetic progressions $\{9k+5\}_{k\in\mathbb{N}}$ and $\{9k+8\}_{k\in\mathbb{N}}$ contains a sum of two triangular numbers. Further, Theorem 1 together with the Prime Number Theorem for arithmetic progressions (Corollary 11.20 in [7]) imply the following.
Corollary 4.
For each pair of integers $r,s>2$, at least one of which is not equal to $4$ and for every pair of $a,b\in\mathbb{N}$, the double sequence
$\{aT(r,m)+bT(s,n)\}_{m,n\in\mathbb{N}}$
misses infinitely many primes from every admissible arithmetic progression.
In other words, one can only describe the primes of the form $am^{2}+bn^{2}$ by a finite set of congruences for certain $a,b\in\mathbb{N}$ (see [3]). However, for any pair of $a,b\in\mathbb{N}$, if we replace either $m^{2}$ or $n^{2}$ above by another polygonal number, the resulting linear combination of polygonal numbers fails to produce infinitely many primes contained in any given admissible arithmetic progression.
Theorem 1 along with Chebyshev’s theorem (Corollary 2.6 in [7]) and Green-Tao Theorem [4] implies:
Corollary 5.
For each pair of coprime positive integers $a$ and $b$, there are arbitrariliy long arithmetic progressions among the primes of the form
$am^{2}+bn^{2}.$
It also follows from Theorem 1 and Chebyshev’s theorem that the sequence of the primes, which are linear combinations (with coprime positive integer coefficients) of two polygonal numbers at least one of which is not a square, has zero relative density with respect to the sequence of all primes. So, possible existence of arbitrary long arithmetic progressions among such primes (see the following table) is undetectable by Green-Tao Theorem [4].
Table 1. Examples of arithmetic progressions among the primes which are sums of two triangular numbers
Length Arithmetic progressions
3 31, 37, 43
4 43, 61, 79, 97
5 61, 151, 241, 331, 421
6 7, 37, 67, 97, 127, 157
7 157, 37747, 75337, 112927, 150517, 188107, 225697
8 3307, 69457, 135607, 201757, 267907, 334057, 400207, 466357
9 3823, 6133, 8443, 10753, 13063, 15373, 17683, 19993, 22303
10 73091, 135461, 197831, 260201, 322571, 384941, 447311, 509681, 572051, 634421
The above table suggests the existence of arbitrarily long arithmetic progressions among the primes which are sums of two triangular number. More generally, we ask whether there are arbitrarily long arithmetic progressions among the primes which are linear combinations of two polygonal numbers with coprime positive integer coefficients.
@@
For each choice of $a,b,r,s\in\mathbb{N}$ with $r,s>2$ and $\gcd(a,b)=1$, are there are arbitrarily long arithmetic progressions among the primes of the form $aT(r,m)+bT(s,n)?$
2. Proofs of the theorems
@@
A Let $P(x,y)=ax^{2}+bxy+cy^{2}+ex+fy+g$ be of an irreducible polynomial in $\mathbb{Z}[x,y]$ such that $\frac{\partial{P}}{\partial x}$ and $\frac{\partial{P}}{\partial y}$ are linearly independent and $P(x,y)$ represents arbitrarily large odd numbers for $x,y\in\mathbb{N}$. Then
@@
Since for all $i\in\{1,2,3\}$, the coefficients of $x^{2}$ and $y^{2}$ in $P_{i}(x,y)$ are positive and even, whereas the coefficient of $xy$ is zero, it follows that if any of the remaining coefficients of $P_{i}$ are odd, then $P_{i}(x,y)$ produces infinitely many odd numbers for $x,y\in\mathbb{N}$. In particular, if the greatest common divisors of the coefficients of $P_{1}$, $P_{2}$, $P_{3}$ are 1, then each of these polynomials represents arbitrarily large odd numbers. In the following we compute the gcd $G$ of the coefficients of $P_{1}$, $P_{2}$ and $P_{3}$. For $i\in\{1,2,3\}$ and $N\in\mathbb{N}$, let
$\pi_{i}(N):=\sum_{\begin{subarray}{c}\text{prime \ }p\hskip 1.0pt\leq N,\\ p\hskip 1.0pt=P_{i}(x,y)\\ \text{for $x,y\in\mathbb{N}$}\end{subarray}}1.$
@@
$\pi_{1}(N)\asymp\begin{cases}\dfrac{N}{(\log N)^{3/2}}&\text{if $\{\mu r,\nu s\}\nsubseteq 2\mathbb{Z}$}\\ \\ 1&\text{otherwise. }\end{cases}$
@@
$\pi_{2}(N)\asymp\begin{cases}\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\nu\notin 2\mathbb{Z}$}\\ \\ \dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ 1&\text{otherwise. }\end{cases}$
@@
$\pi_{3}(N)\asymp\begin{cases}\dfrac{N}{\log N}&\text{if }r=s=4\text{ and $\mu+\nu\notin 2\mathbb{Z}$}\\ \\ \dfrac{N}{(\log N)^{3/2}}&\text{if $\{r,s\}\neq\{4\}$ and $\{r,s,\mu+\nu\}\nsubseteq 2\mathbb{Z}$}\\ \\ 1&\text{otherwise. }\end{cases}$
@@
[…]
Acknowledgments
All the computations for Table 1 was done using PARI/GP. Also, Bhattacharya thankfully acknowledges the support from the Department of Science and Technology, Govt. of India through the INSPIRE Faculty Award IFA17-MA104 and Rahaman thankfully acknowledges the support from the Ministry of Education, Govt. of India through the Prime Minister Research Fellowship 0501972.